OpenStudy (anonymous):

I am having trouble with this problem. I cannot get the answer in the back of the book. y'=(x+y)/(2x)

OpenStudy (anonymous):

solve the homogeneous equation

OpenStudy (anonymous):

nonlinear

OpenStudy (anonymous):

you can rewrite the equation in this way \[y'=\frac{ x(1+\frac{ y }{ x })0 }{ 2x }=\frac{ 1+\frac{ y }{ x } }{ 2 }\]

OpenStudy (anonymous):

without the zero there , missclicked

OpenStudy (anonymous):

now u can set \[u=\frac{ y }{ x }=>y=ux=>y'=u'x+u\]

OpenStudy (anonymous):

so now just substitude for y' and for u

OpenStudy (anonymous):

that what I have been doing and I get \[\frac{ (x-y)^{2} }{ (x)^{3} }=c\]

OpenStudy (anonymous):

hm i assume u have some mistake

OpenStudy (anonymous):

I do not know what I did wrong. I keep getting the same answer.

OpenStudy (anonymous):

ok i tell u what i get , i get smt like \[x-constant*\sqrt{x}=y\]

OpenStudy (anonymous):

did you get to the point where u should find integrals of this \[\int\limits_{}^{}{\frac{ dx }{ x }}=\int\limits_{}^{}{\frac{ 2du }{ 1-u }}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so what did u get for the integrals?

OpenStudy (anonymous):

\[\ln x +C=2\ln (1-v)\]

OpenStudy (anonymous):

ok seems like you forgot a "-" sign in front of 2ln(1-u)

OpenStudy (anonymous):

oh that it. this problem has been troubling me for hours. Such I stupid Mistake. Thanks a lot!

OpenStudy (anonymous):

did u get right answer now ? :D and yeah such a shame losing so much time at just a - sign xD

OpenStudy (anonymous):

yea I got the right answer.

OpenStudy (anonymous):

cute , good luck ^^